# DPZV: Resource Efficient ZO Optimization For Differentially Private VFL

### v1

@@ 7.2 DPZV Comparison to Baselines
Figure 2 presents the performance evaluation of DPZV against all baselines on image classification tasks without DP constraints. To ensure a fair comparison, model delay has been manually adjusted for all methods based on the per-batch computation time on a single client. The results show that DPZV consistently outperforms all baselines on MNIST and CIFAR-10. On ModelNet40, where training is more challenging due to larger models and a greater number of clients, DPZV maintains a competitive advantage over other ZO-based methods while achieving accuracy comparable to the first-order baseline VAFL. The clipping operation in DPZV effectively controls gradient magnitudes, mitigating the instability caused by noisy ZO gradient estimators. This contributes to improved training stability. Furthermore, the memory-efficient ZO optimization significantly enhances computational efficiency, leading to a faster convergence rate compared to first-order methods.
Communication Cost: Table 1 shows the communication cost of one client for all algorithms performing forward (client-server) communication and backward (server-client) communication, where the communication cost relies on the dimension of the embeddings and batch size. In MNIST, we used an embedding size of $60$, while in CIFAR-10 and ModelNet40, the embedding size is set at $128$. Here, we assume an embedding size of $128$ for comparison. We can observe from that DPZV significantly reduces the backward communication cost from the server to the client compared to VAFL and VFL-CZOFO, as ZO optimization eliminates the need for backward propagation. Compare to ZOO-VFL, DPZV is able to reduce the forwarding communication by leveraging the asymmetric update mechanism. Our method reduces the total communication cost by nearly half of all existing methods. This demonstrated the superiority of our method to perform communication efficient training. Figure 3 evaluates the performance of VFL methods under the same communication cost. We observe that DPZV achieve a bigger lead compared to all baselines under the same communication cost, reaching the same level of convergence at a much smaller communication cost.
Memory Cost: Figure 4 compares the GPU memory consumption, with memory values normalized for readability. We compare the memory cost on larger models, where we use ResNet for image classification and BERT for sequence classification. We record the highest memory peak to show the total required memory for each method on training large models. We observe that DPZV requires memory approximately equal to the model size, whereas the first-order method VAFL demands more than twice the model size. Compared to ZOO-VFL, DPZV achieves further memory savings by leveraging MeZO. These results highlight the scalability of DPZV, making it well-suited for deploying large pretrained language models in VFL scenarios.
7.3 Impact of different DP levels
We further compare the utility of DPZV with baselines under various DP levels. VAFL achieves DP by adding random noise to the output of each local embedding. VFL-CZOFO utilize the intrinsic privacy protection of ZO optimization to give DP guarantee. However, the intrinsic DP level sometimes falls below the required level and is not adjustable. ZOO-VFL does not give DP guarantee. To enforce DP and compare all methods fairly, we adopt the basic DP-SGD [32] method on VFL-CZOFO and ZOO-VFL by clipping gradient and adding noise.
In Table 2 we evaluate three levels of DP guarantees: (1) no DP noise injected, (2) noise injected to ensure $\epsilon=10$, and (3) noise injected to ensure $\epsilon=1$. As $\epsilon$ decreases, the performance of VAFL and ZOO-VFL deteriorates significantly. While VFL-CZOFO exhibits a relatively smaller accuracy drop compared to VAFL and ZOO-VFL, the reduction remains substantial, hindering the system from achieving an optimal model. In contrast, our DPZV maintains consistently high performance even under strict privacy budgets ($\epsilon\leq 1$), with only minor deviations from its No DP accuracy. These results show the advantage of DPZV, where memory-efficient zeroth-order updates significantly enhance the training stability of DP-protected systems.
8 Conclusion
@@
In this work, we propose DPZV, the first zero-order VFL framework that achieves low communication overhead, reduced computational cost, and robust privacy guarantees simultaneously. Through rigorous theoretical analysis, we establish the strong convergence properties of our algorithm. Extensive experiments demonstrate that DPZV outperforms all baselines across various privacy levels and cost evaluations. Our method also features a lightweight memory footprint, enabling enhanced performance on large language models. These results provide valuable insights for future advancements in VFL and distributed ZO optimization methods.
@@ 2. $(ii)$
$\leq$ $\frac{1}{\alpha T}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{1}{\alpha}\sum_{m=0}^{M}q_{m}\eta_{m}\frac{L^{2}}{8}\lambda^{2}d_{m}^{3}+\frac{1}{\alpha}\sum_{m=0}^{M}q_{m}\eta_{m}^{2}L\left(\frac{4}{B}d_{m}\sigma_{s}^{2}+\frac{(4-B)L^{2}}{4B}\lambda^{2}d_{m}^{3}+\frac{2P}{B}C^{2}d_{m}+2\sigma_{dp}^{2}d_{m}\right)$
@@ Appendix D Intermediate Lemmas
$=$ $F_{\lambda}({\bm{w}}^{t})-\eta_{0}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\rangle+\frac{L\eta_{0}^{2}}{2}\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}-\eta_{m}\langle\nabla_{w_{m}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{m}(\tilde{{\bm{w}}}^{t})\rangle+\frac{L\eta_{m}^{2}}{2}\mbox{$\left\lVert\breve{G}^{t}_{m}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}$
$\mathbb{E}\left[F_{\lambda}({\bm{w}}^{t+1})\right]\leq$ $\mathbb{E}\left[F_{\lambda}({\bm{w}}^{t})\right]\underbrace{-\eta_{0}\mathbb{E}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\rangle}_{{\mathcal{E}}_{1}}+\underbrace{\frac{L\eta_{0}^{2}}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]}_{{\mathcal{E}}_{2}}\underbrace{-\eta_{m_{k}}\mathbb{E}\langle\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\rangle}_{{\mathcal{E}}_{3}}$
$+\underbrace{\frac{L\eta_{m_{k}}^{2}}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]}_{{\mathcal{E}}_{4}}$
@@
$=$ $\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})+(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
$\overset{2)}{\leq}$ $2\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]+2(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
@@
$\frac{1}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]\leq$ $\frac{4(1-P)d_{m}}{B}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]+2L^{2}\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]+{\mathcal{G}}_{m}$ (31)
@@
$+$ $\frac{4(1-P)d_{m_{k}}L\eta_{m_{k}}^{2}}{B}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)L\eta_{m_{k}}^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]+2L^{3}\eta_{m_{k}}^{2}\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]+L\eta_{m_{k}}^{2}{\mathcal{G}}_{m}$
@@
-- $\sum_{m=1}^{M}q_{m}\eta_{m}(1-P)\left(\frac{1}{2}-\eta_{m}L\right)\left(\frac{1}{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]-\frac{L^{2}}{4}\lambda^{2}d_{m}^{3}\right)+\sum_{m=1}^{M}q_{m}\eta_{m}L\left(\frac{1}{2}+2\eta_{m}L^{2}\right)\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]$
@@ E.2 Protection Against Post-Training Attacks
Appendix F Choice of all Hyperparameters

### v2

@@ 7 Experiments
Comparison with Baselines. Figure 2 presents the performance evaluation of DPZV against all baselines on image classification tasks under certain DP levels. The results show that DPZV consistently outperforms all baselines on MNIST and CIFAR-10 under strict privacy budget ($\epsilon=1,\delta=1\times 10^{-3}$). On ModelNet40, where training is more challenging due to larger models and a greater number of clients, we slightly relax the DP constraint ($\epsilon=10$), and observe that DPZV maintains a competitive advantage over other ZO-based methods while achieving accuracy comparable to the first-order baseline VAFL. The scalar noise injected in DPZV effectively constrains the noise magnitude, mitigating the instability typically introduced by noisy zeroth-order gradient estimators. This leads to more stable training dynamics than other ZO based methods. Moreover, in contrast to the high-dimensional vector noise used in first-order baseline VAFL, scalar noise incurs significantly less loss, resulting in superior performance under stringent privacy budgets.
Privacy-Accuracy Tradeoff. Figure 3 presents the accuracy-privacy tradeoff across four benchmark datasets: MNIST, CIFAR-10, ModelNet40, and Amazon Reviews. We evaluate the robustness of our proposed method DPZV under various privacy budgets $\epsilon$, while fixing the failure probability $\delta=1\times 10^{-3}$. Across all tasks, DPZV consistently achieves the highest accuracy under tight privacy regimes ($\epsilon\leq 1$), indicating its ability to maintain model utility even with stringent differential privacy constraints. Notably, on MNIST and ModelNet40, DPZV shows only a marginal drop in accuracy as $\epsilon$ decreases, while the competing methods suffer substantial degradation. For instance, on CIFAR-10 at $\epsilon=0.1$, DPZV maintains over 90% accuracy, whereas ZOFO and ZOO-VFL fall below 30%. Similar trends are observed on Amazon Reviews, where DPZV consistently outperforms baselines under strict privacy, especially at $\epsilon=1$ and $\epsilon=5$. These results validate the effectiveness of our scalar-noise-based DP mechanism in balancing utility and privacy, and highlight its advantage in real-world privacy-sensitive federated learning scenarios.
Figure 4: Effect of gradient clipping on DPZV under non-private settings. We compare models trained with and without clipping of the ZO information. Across both MNIST and CIFAR-10, clipping accelerates convergence and stabilizes training.
Figure 5: Effect of clipping threshold on DPZV under differential privacy constraint $\epsilon=1$. While all clipping levels perform similarly on MNIST, smaller thresholds ($C=10$) significantly improve accuracy and convergence on CIFAR-10.
Effect of Clipping in Non-Private Scenarios. We observe in Figure 3 that even under no privacy constraints ($\epsilon=\infty$), DPZV outperforms other ZO based algorithms. A key distinction lies in DPZV’s use of scalar clipping on ZO information, while originally introduced for differential privacy, also acts as a form of gradient regularization. This regularization effect has been shown to improve convergence in prior work [37], and we observe similar benefits here. Figure 5 verifies our insight by comparing the convergence behavior of DPZV with and without gradient clipping under a non-private setting. The plots show test loss versus communication rounds. In both datasets, applying clipping to the ZO information significantly improves convergence speed. For MNIST, clipped DPZV reaches low test loss much faster and stabilizes more smoothly. On CIFAR-10, the clipped version also demonstrates consistently lower loss throughout training. These results suggest that clipping not only stabilizes training but also enhances convergence efficiency, even when privacy is not enforced.
Impact of Sensitivity Level. Figure 5 presents the performance of DPZV under different sensitivity levels, which is controlled by the clipping threshold $C$. We apply a fixed privacy budget of $\epsilon=1,\delta=1\times 10^{-3}$ on both MNIST and CIFAR-10. In the MNIST setting, all clipping levels achieve similar convergence and final accuracy, suggesting the model is robust to the choice of $C$ in simple tasks. In contrast, the CIFAR-10 results highlight a pronounced impact: smaller clipping values (e.g., $C=10$) yield better accuracy and stability over training. Larger thresholds, such as $C=500$ degrade performance, likely due to the excessive noise required to satisfy DP constraints. These results emphasize the importance of careful clipping calibration in more complex settings to ensure a good privacy-utility tradeoff.
8 Conclusion and Limitations
@@
In this work, we propose DPZV, the first zero-order VFL framework that achieves robust and controllable privacy guarantees while maintaining strong model utility. Through rigorous theoretical analysis, we establish the strong convergence properties of our algorithm. Extensive experiments demonstrate that DPZV outperforms all baselines across various privacy levels and cost evaluations. Our method also features a lightweight memory footprint, enabling enhanced performance on large language models. These results provide valuable insights for future advancements in VFL and distributed ZO optimization methods.
We identify several potential limitations of our work. Notably, the convergence rate established in our theoretical analysis remains dimension-dependent, consistent with prior results in centralized ZO optimization. Extending the dimension-free convergence guarantees shown in [26] to the VFL setting poses a nontrivial challenge due to the inherently distributed and feature-partitioned nature of VFL. Investigating such extensions is a promising direction for future work.
Appendix
[…]
@@ 2. $(ii)$
$\leq$ $\frac{1}{\alpha T}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{1}{\alpha}\sum_{m=0}^{M}q_{m}\eta_{m}\frac{L^{2}}{8}\lambda^{2}d_{m}^{3}$
$+\frac{1}{\alpha}\sum_{m=0}^{M}q_{m}\eta_{m}^{2}L\left(\frac{4}{B}d_{m}\sigma_{s}^{2}+\frac{(4-B)L^{2}}{4B}\lambda^{2}d_{m}^{3}+\frac{2P}{B}C^{2}d_{m}+2\sigma_{dp}^{2}d_{m}\right)$
@@ Appendix D Intermediate Lemmas
$=$ $F_{\lambda}({\bm{w}}^{t})-\eta_{0}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\rangle+\frac{L\eta_{0}^{2}}{2}\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}$
$-\eta_{m}\langle\nabla_{w_{m}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{m}(\tilde{{\bm{w}}}^{t})\rangle+\frac{L\eta_{m}^{2}}{2}\mbox{$\left\lVert\breve{G}^{t}_{m}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}$
$\mathbb{E}\left[F_{\lambda}({\bm{w}}^{t+1})\right]\leq$ $\mathbb{E}\left[F_{\lambda}({\bm{w}}^{t})\right]\underbrace{-\eta_{0}\mathbb{E}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\rangle}_{{\mathcal{E}}_{1}}+\underbrace{\frac{L\eta_{0}^{2}}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]}_{{\mathcal{E}}_{2}}$
$\underbrace{-\eta_{m_{k}}\mathbb{E}\langle\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t}),\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\rangle}_{{\mathcal{E}}_{3}}+\underbrace{\frac{L\eta_{m_{k}}^{2}}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]}_{{\mathcal{E}}_{4}}$
@@
$=$ $\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})+(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
$+(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
$\overset{2)}{\leq}$ $2\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]+2(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
$+(1-P)^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
@@
$\frac{1}{2}\mathbb{E}\left[\mbox{$\left\lVert\breve{G}^{t}_{m_{k}}(\tilde{{\bm{w}}}^{t})\right\rVert$}^{2}\right]\leq$ $\frac{4(1-P)d_{m}}{B}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$ (30)
$+2L^{2}\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]+{\mathcal{G}}_{m}$ (31)
@@
$+$ $\frac{4(1-P)d_{m_{k}}L\eta_{m_{k}}^{2}}{B}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]+(1-P)L\eta_{m_{k}}^{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m_{k}}}F_{\lambda}({\bm{w}}^{t})\right\rVert$}^{2}\right]$
$+$ $2L^{3}\eta_{m_{k}}^{2}\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]+L\eta_{m_{k}}^{2}{\mathcal{G}}_{m}$
@@
-- $\sum_{m=1}^{M}q_{m}\eta_{m}(1-P)\left(\frac{1}{2}-\eta_{m}L\right)\left(\frac{1}{2}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{w_{m}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]-\frac{L^{2}}{4}\lambda^{2}d_{m}^{3}\right)$
$+$ $\sum_{m=1}^{M}q_{m}\eta_{m}L\left(\frac{1}{2}+2\eta_{m}L^{2}\right)\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]$
@@ E.2 Protection Against Post-Training Attacks
Appendix F Memory Cost
Figure 6 compares the GPU memory consumption, with memory values normalized for readability. We compare the memory cost on larger models, where we use ResNet for image classification and BERT for sequence classification. We record the highest memory peak to show the total required memory for each method on training large models. We observe that DPZV requires memory approximately equal to the model size, whereas the first-order method VAFL demands more than twice the model size. Compared to ZOO-VFL, DPZV achieves further memory savings by leveraging MeZO. These results highlight the scalability of DPZV, making it well-suited for deploying large pretrained language models in VFL scenarios.
Appendix G Choice of all Hyperparameters